OpenStudy (anonymous):

One gallon of paint (volume = 3.78 10-3 m3) covers an area of 20.0 m2. What is the thickness of the fresh paint on the wall?

OpenStudy (anonymous):

Well... just try thinking of it as a cuboid of paint with the volume 3.78 dm³ and a "base area" (don't know if that's the right term in english) of 20m²...

OpenStudy (anonymous):

i have no idea lol..thats just the question given to me =(

OpenStudy (anonymous):

you know how to calculate the volume of a cuboid?

OpenStudy (anonymous):

i dont..we didnt even learn it and its on my hw

OpenStudy (anonymous):

Well, it's just V = a*b*c or Volume = depth*length*height

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

And I called A = depth*lenght the "base area"... then the height of the cuboid would be the thickness of the paint.. Do you understand that ahm thought? :P

OpenStudy (anonymous):

i get the formula but i dont get how to do it with the Volume being written like that

OpenStudy (anonymous):

but you get how you can think the layer of paint as a very flat/shallow cuboid? The rectangular base area is A = 20m² - to get the volume, we have to multiply this by the height..\[V = A * h\]but we already know the volume (3.78 dm³) and want to get the height (thickness) instead... so we ahm.. change the formula to\[h = {V \over A} \]Now you just have to plug in the values ;)